Seether, Inc., has the following two mutually exclusive projects available.

Year Project R Project S
0 –$ 55,000      –$ 76,000     
1 21,000      20,000     
2 22,000      20,000     
3 19,000      35,000     
4 12,000      30,000     
5 9,000      10,000     
Requirement 1:

What is the crossover rate for these two projects? (Do not round intermediate calculations. Enter your answer as a percentage rounded to 2 decimal places (e.g., 32.16).)

  Internal rate of return %
Requirement 2:

What is the NPV of each project at the crossover rate? (Do not round intermediate calculations. Round your answers to 2 decimal places (e.g., 32.16).)

    NPV
  Project R $   
  Project S $   
Answer #1

Answer

Requirement -1 : Crossover Rate of two projects is the rate of return where the NPV of two projects becomes equal.

NPVR = NPVS

-55000 + 21000/ (1+r) +22000/ (1+r)2 + 19000/ (1+r)3 + 12000/ (1+r)4 + 9000/ (1+r)5 = -76000 + 20000/ (1+r) +20000/ (1+r)2 + 35000/ (1+r)3 + 30000/ (1+r)4 + 10000/ (1+r)5

21000 + 1000/ (1+r) +2000/ (1+r)2 - 16000/ (1+r)3 - 18000/ (1+r)4 - 1000/ (1+r)5 = 0

r= Cross over Rate = 11.87 %

IRRR = 18.26 %

IRRS = 15.92 %

Requirement -2 :

PV at interest rate of 11.87%

NPVR = NPVS ( At Cross Over Rate) = $ 7720
Answer #2

1.
NPV of Project R=-55000+21000/(1+r)+22000/(1+r)^2+19000/(1+r)^3+12000/(1+r)^4+9000/(1+r)^5=7721.00678391295

2.
NPV of Project S=-76000+20000/(1+r)+20000/(1+r)^2+35000/(1+r)^3+30000/(1+r)^4+10000/(1+r)^5=7721.00678400605

3.
Crossover rate:
-55000+21000/(1+r)+22000/(1+r)^2+19000/(1+r)^3+12000/(1+r)^4+9000/(1+r)^5=-76000+20000/(1+r)+20000/(1+r)^2+35000/(1+r)^3+30000/(1+r)^4+10000/(1+r)^5

=>r=11.87%